Operation Summary for September 18, 2009 - September 21, 2009 |
The Integrated Luminosity 9/14/09 to 9/20/09 was 19.01 pb-1. Friday Sept. 18 The day shift began with Operations monitoring the Tevatron (TeV) store 7188 (15.18E30) and the Recycler stash (275.9E10), with the Antiproton Source stacking (25.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the MiniBooNE and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. Operations terminated store 7188 at 9:02 AM. Operations put the TeV into shot setup and established store 7189 at 11:40 AM with an average initial luminosity (AIL) of 175.61E30. TeV personnel replaced the sector D1 BPM crate at 3:14 PM. Operations rebooted a Linac low-level RF station (LRF3) VME at 3:20 PM. Operations terminated store 7189 at 9:31 PM. Operations established store 7190 at 10:58 PM with an AIL of 180.79E30. There were several Booster and Linac RF station trips during the evening shift that caused downtime. Saturday Sept. 19 The midnight shift began with Operations monitoring store 7190 (144.03E30) and the stash (112.3E10), with Pbar stacking (23.2E10), and with NuMI and MiniBooNE taking beam. At 6:55 AM, a MCR console failed while operators were beginning a Recycler stash. Operators could not reset the transfer sequencer. Operations contacted controls personnel who helped restart the console and sequencer. There were several Booster and Linac RF station trips during the midnight shift that caused downtime. At 10:08 AM, the MI-10 lower power supply interlock loop (PSLOOP) tripped and opened the VCB. The VCB would not close, nor would the house reset. Operations called in EE Support personnel to investigate. MI resumed ramping at 5:07 PM. EE Support personnel reported at 5:28 PM that they found and replaced a bad PLC board. Operations terminated store 7190 at 6:09 PM. Operations established store 7191 at 7:54 PM with an AIL of 192.98E30. At 10:12 PM, TeV sectors A1 and F4 quenched. Store 7191 was lost. Operations reported later that the quench was due to a kicker (PAK3) pre-fire. Sunday Sept. 20 The midnight shift began with the TeV recovering from the quench, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. At 12:54 AM, a Linac RF station (LRF3) tripped off for the first of several times. It reset. Operations put the TeV through a dry squeeze at 1:40 AM. At 1:58 AM, LRF3 tripped off several more times. It reset. Operations put the TeV through a wet squeeze at 2:21 AM and then again at 3:06 AM. Operations put the TeV into shot setup at 3:43 AM. Operations established store 7194 at 5:44 AM with an AIL of 171.62E30. Operations reported other short down times throughout the midnight shift due to Klystron RF stations (KRF1 & KRF4A), Linac quadrupole power supplies (QPS104, 132, & 210), BRF5 & 8, and the Booster East Anode Power Supply (EAPS). Operations terminated store 7194 at 7 PM. Operations put the TeV into shot setup at 7:11 PM. At 7:36 PM, the TeV locked at 90 GeV. Operators investigated and found a problem with the TeV sector D4 cryo system lead temperature. Operations contacted the Cryo Coordinator. At 8:05 PM, Operations reported a TeV sector C1 HFU capacitor bank had failed. Operators began replacing the bank. Operators had the C1 HFU capacitor bank replaced by 8:56 PM. Operations put the TeV into shot setup and established store 7196 at 11:54 PM with an AIL of 220.38E30. Monday Sept. 21 The midnight shift began with Operations monitoring store 7196 and the stack, with Pbar stacking, and with NuMI and MiniBooNE taking beam. At 5:04 AM, LRF3 tripped off for the first of two times in 12 minutes. It reset. Operations turned off the MiniBooNE beam at 7:43 AM for scheduled HVAC work. The Plans for Monday The plans for today are to stack, stash, and store. Switchyard personnel will take beam to their dump on Tuesday. |
|
More Information Comments and Suggestions |
| last modified 09/21/2009 email Fermilab |
| Security, Privacy, Legal | |